OpenStudy (anonymous):

Solve for (g-7f)(8) when f(x) = 4x^2-5x+6

jimthompson5910 (jim_thompson5910):

what is g(x)?

OpenStudy (anonymous):

Sorry it's g(x)=x^2+4

jimthompson5910 (jim_thompson5910):

(g-7f)(8) is the same as g(8) - 7*f(8)

jimthompson5910 (jim_thompson5910):

so you first need to find f(8) and g(8)

jimthompson5910 (jim_thompson5910):

f(x) = 4x^2-5x+6 f(8) = 4(8)^2-5(8)+6 f(8) = ??